Consider the following inverse demand of a market: P = 100 – Q. Suppose the firms...

  1. Consider the following inverse demand of a market: P = 100 – Q. Suppose the firms in the market are identical and have MC = 80.

    1. If a process innovation reduces MC to 70, is the innovation drastic or non-drastic and

      why?

    2. If a process innovation reduces MC to 50, is the innovation drastic or non-drastic and

      why?

    3. What is the threshold MC, below which the innovation becomes drastic innovation?
Answer #1

If a innovation makes firm a monopolist in the market then innovation is drastic , otherwise it is non drastic.

If MR=100-2Q

Innovation reduces MC=70

MR=MC

100-2Q=70

Q=15

P=100-15=85

innovator cannot charge $85 because rivals have unit cost of $80 and could under price.firm doesn't get monopoly power.

So innovation is non drastic.

Innovation reduces MC=50

MR=MC

100-2Q=50

Q=25

P=100-25=75

Innovator can charge optimal monopoly price ($75) and still undercut rivals whose unit cost is 80$.and firm gets monopoly power.

So innovation is drastic.

P=80

Q=100-80=20

MR=100-2Q=

MR=100-2*20=60

So threshold MC=60 ,below which innovation becomes drastic innovation